# High Fidelity Scene Text Synthesis

### v1

@@ 4.2 Results
4.3 Ablation Studies and Effectiveness Analysis
Ablation on Losses: In our ablation study on key design losses within our heuristic alternate optimization pipeline, we observe a significant drop in performance across all metrics when these losses are removed, as shown in Tab. 4.3. Additionally, these results highlight that the masked diffusion loss and cross-attention loss alone are insufficient for achieving satisfactory character representation. This may be attributed to the inherent noise in the latent character masks. To sum up, these ablation results underscore the effectiveness of each objective within our heuristic alternate optimization strategy in improving various aspects of synthesized scene texts.
Choice of $\alpha$ and $\beta$: In our experiments, we vary the values of $\alpha$ and $\beta$ within the range [0.001, 0.01, 0.1, 1]. The results are presented in Tab. 4.3. Initially, we focus on exploring the impact of $\alpha$ while setting $\beta$ = 0. We find that the optimal performance in text reconstruction occurs when $\alpha$ = 0.001. However, this choice yields poor results in editing tasks. Consequently, we opt for a more balanced approach, selecting $\alpha$ = 0.01. Subsequently, with $\alpha$ fixed at 0.01, we investigate the effect of $\beta$. We find that the optimal performance is achieved when $\beta$ = 0.001.
@@ 4.3 Ablation Studies and Effectiveness Analysis
Table 2: Ablation results on different settings.
@@
Recon Editing FID
Base 0.362 0.217 51.29
+$\mathcal{L}_{mask}$ 0.621 0.559 45.21
+$\mathcal{L}_{attn}$ 0.698 0.623 36.98
+$\mathcal{L}_{align}$ 0.873 0.782 26.53
+$\mathcal{L}_{id}$ 0.940 0.887 15.79
@@ 5 Conclusion
Appendix
Appendix A Datasets
Several datasets are used in our experiments: (1) SynthText [19] comprises 800,000 images containing around 8 million synthetic word instances. Each text instance is annotated with its corresponding text string, along with word-level and character-level bounding boxes. (2) LAION-OCR [7] contains 9,194,613 filtered high-quality text images including advertisements, notes, posters, covers, memes, logos, etc. (3) ICDAR13 [20] is widely recognized as the benchmark dataset for evaluating near-horizontal text detection, which consists of 233 test images used for evaluation purposes. (4) TextSeg [21] comprises 4,024 real-world text images sourced from various sources such as posters, greeting cards, book covers, logos, road signs, billboards, digital designs, handwritten notes, and more. Note that only SynthText and LAION-OCR datasets provide the character segmentation maps. We train our model using the training subsets of these datasets and randomly select 100 images from test subsets for testing.
Appendix B Training Configurations
We train our model based on the pre-trained checkpoint of Stable Diffusion (v2.0) and text encoder in [10]. The pre-trained Vision Transformer [25] is used as our image encoder. The model is fine-tuned using 4 NVIDIA A100 GPUs on LAION-OCR for 200k steps, SynthText for 150k steps, TextSeg for 50k steps, and ICDAR13 for an additional 10k steps. During training, we set $\alpha$ to 0.01, and $\beta$ to 0.001. Additionally, we utilize a batch size of 16 and a learning rate of 5 $\times$ $10^{-5}$.
Appendix C Experimental Settings for Baseline Methods
In this study, we conduct a comprehensive comparison of our method against several state-of-the-art baselines, encompassing both GAN-based and diffusion-based approaches. Specifically, we evaluate our method against MOSTEL [1], Stable Diffusion-inpainting (v2.0) [17], DiffSTE [6], TextDiffuser [7], and UDiffText [10]. For MOSTEL, we utilize it to generate text within the masked region and then integrate the output back into the original image. Regarding Stable Diffusion, we employ the publicly available pre-trained model "stable-diffusion-2-inpainting" from Hugging Face, setting its prompt as "[word to be rendered]" for a fair comparison. For TextDiffuser, we utilize their inpainting variant, where the desired text is rendered in a standard font (Arial) within the masked region, serving as input for their proposed segmentor. Finally, for UDiffText and DiffSTE, we follow the settings outlined in their respective original papers.
Table 4: Ablation study results on balanced supervision.
Setting Average SeqAcc
Recon Editing FID mIoU
unsupervised 0.212 0.157 62.36 0.203
supervised 0.862 0.813 14.92 0.617
balanced supervision (Ours) 0.940 0.887 12.13 0.722
Appendix D Ablation Study on Balanced Supervision
We analyze the efficacy of our balanced supervision for character attention by comparing it with unsupervised and self-supervised learning approaches. We additionally incorporate the mean Intersection over Union (mIoU) metric to evaluate the alignment between their latent character masks and ground truth character segmentation masks on the LAION-OCR dataset. As illustrated in Tab. 4, when conducting unsupervised learning, we observed a significant deterioration in the model’s performance across all metrics. This outcome is reasonable since the model lacks prior knowledge about determining the ideal character locations at the outset. Consequently, it fails to concentrate attention effectively around character regions, leading to significant deviations between the characters’ positions in generated latent character masks and their actual positions. As a result, the autonomous alternate optimization process is adversely affected.
Furthermore, we explore self-supervised learning by guiding attention calibration using the cross-entropy objective between latent character masks and character segmentation masks. Although self-supervised learning demonstrates a notable improvement compared to unsupervised learning, there remains a gap when compared to our balanced supervision. This discrepancy arises from the overly strong constraint of the characters’ position, limiting the model’s flexibility in estimating optimal positions, which hinders its ability to adapt to varied and complex scenarios.
Appendix E Limitations
While our method demonstrates promising capabilities in synthesizing scene text, it is limited by its inability to simultaneously modify multiple regions within an image, which restricts its applicability. Future research will explore techniques to address this limitation, aiming to develop more efficient and versatile text synthesis methods capable of simultaneously generating multiple texts within an image. Besides, its application prompts considerations regarding privacy. The generation of realistic text, including personal signatures or identifiable information, may pose risks if misused, potentially compromising individuals’ privacy and security. These concerns underscore the importance of implementing robust safeguards and ethical guidelines to address potential privacy risks and ensure the responsible use of this technology.
Appendix F Societal Impacts
The advancement of scene text synthesis technology in our work holds significant societal implications. Firstly, it contributes to cultural preservation by enabling the generation of text in diverse styles and languages, aiding in the digitization and conservation of historical scripts and manuscripts. Additionally, our method has applications in art, design, and advertising, empowering creators to produce visually captivating compositions and typography designs. However, ethical considerations surrounding the potential misuse of synthesized text for fraudulent purposes must be carefully addressed through the development of robust safeguards and guidelines. Therefore, while our work offers promising possibilities for text synthesis, it necessitates thoughtful consideration of its societal impacts and ethical implications.
Appendix G Additional Visualized Attention Results
Additional visualized attention cases of all characters across several steps during training are exhibited in Fig. 9.
Appendix H Additional Qualitative Comparative Results
Additional qualitative comparisons against the baselines are exhibited in Fig. 10.
Appendix I Additional Visual Results
Additional visual results generated by our DreamText are exhibited in Fig. 11 and Fig. 12.
Figure 9: Additional visualized attention cases of all characters across several steps during training.
Figure 10: Additional qualitative comparative results against state-of-the-art methods.
Figure 11: Additional visual results generated by our DreamText.
Figure 12: Additional visual results generated by our DreamText.

### v2

@@ Results
Ablation Studies and Effectiveness Analysis
Ablation on Losses: We employ Stable Diffusion (v2.0) inpainting version as the base method. As shown in Tab. 2, we observe a significant drop in performance across all metrics when these losses are removed. Additionally, these results highlight that the masked diffusion loss and cross-attention loss alone are insufficient for achieving satisfactory character representation. This may be attributed to the inherent noise in the latent character masks. To sum up, these ablation results underscore the effectiveness of each objective within our heuristic alternate optimization strategy in improving various aspects of synthesized scene texts.
Choice of $\alpha$ and $\beta$: In our experiments, we vary the values of $\alpha$ and $\beta$ within the range [0.001, 0.01, 0.1, 1]. The results are presented in Tab. 3. Initially, we focus on exploring the impact of $\alpha$ while setting $\beta$ = 0. We find that the optimal performance in text reconstruction occurs when $\alpha$ = 0.001. However, this choice yields poor results in editing tasks. Consequently, we opt for a more balanced approach, selecting $\alpha$ = 0.01. Subsequently, with $\alpha$ fixed at 0.01, we investigate the effect of $\beta$. We find that the optimal performance is achieved when $\beta$ = 0.001.
@@
Table 2: Accumulated results of losses.
@@
Recon Editing FID LPIPS
Base 0.218 0.060 26.78 0.0696
+$\mathcal{L}_{mask}$ 0.425 0.259 23.21 0.0528
+$\mathcal{L}_{attn}$ 0.698 0.532 19.72 0.0483
+$\mathcal{L}_{align}$ 0.884 0.801 15.41 0.0392
+$\mathcal{L}_{id}$ 0.940 0.887 12.13 0.0328